# Synaptic Pruning: A Biological Inspiration for Deep Learning Regularization

### v1

@@ Abstract
Introduction
Synaptic pruning is a key neuro-developmental process in biological brains that removes weak connections to improve neural efficiency. In artificial neural networks, standard dropout regularization randomly deactivates neurons during training but overlooks the activity-dependent nature of biological pruning. This study proposes a novel regularization method that better reflects biological pruning by gradually eliminating connections based on their contribution to model performance.
Methods
We propose a magnitude-based synaptic pruning approach that integrates directly into the training loop of deep learning models as a replacement for dropout layers. The method computes weight importance from absolute magnitudes across all layers, applying a cubic schedule to gradually increase global sparsity during training. At set intervals, pruning masks are updated by thresholding weights, permanently removing low-importance connections while retaining gradient flow for active weights. This continuous, data-driven pruning eliminates the need for separate pruning and fine-tuning phases.
Results
Experimental results across multiple time series forecasting deep neural network architectures, including Recurrent Neural Networks, Long Short Term Memory, and Patch Time Series Transformer on four different forecasting datasets demonstrate the efficacy of our synaptic pruning method. The method achieved the best overall performance ranking across all architectures, with statistically significant improvements confirmed by Friedman tests (p <0.01) on all datasets. Notably, our proposed method achieves substantial improvements in financial forecasting tasks, reducing Mean Absolute Error by up to 20% compared to models using no dropout or standard dropout for regularization, with reductions reaching up to 52% in select transformer models.
Conclusion
The proposed synaptic pruning mechanism advances regularization by coupling dynamic weight elimination with progressive sparsification during training. Its ease of integration with diverse architectures, coupled with superior forecasting accuracy, underscores its potential as a practical alternative to other dropout techniques, especially in domains like financial time series forecasting, where precision and efficiency are critical.
@@ Conclusion
journal: The Journal of Neuroscience Methods
@@ 1 Introduction
Synaptic pruning is a fundamental neuro-developmental process whereby weak or redundant synaptic connections are selectively eliminated to enhance the efficiency of neuronal networks [1]. This process is highly activity-dependent, meaning that synapses with lower activity or utility are more likely to be pruned, resulting in improved energy efficiency and functional specialization [2]. During critical periods of development, the human brain initially overproduces synapses, creating a dense network of connections that undergoes substantial refinement through adolescence and early adulthood [3, 4]. This refinement is essential for optimal cognitive function and normal brain development, with disruptions in pruning mechanisms associated with various neuro-developmental disorders [5, 6].
In the field of artificial neural networks, researchers have drawn inspiration from biological synaptic pruning to develop methods for improving network efficiency and generalization capacity [7, 8, 9]. Dropout regularization [10] functions by randomly deactivating, or ”dropping out,” a subset of neurons during each training iteration. This process forces the network to avoid over-reliance on any particular neuron, promoting redundancy and robustness in the learned representations. By preventing neurons from co-adapting too strongly to specific patterns in the training data, dropout reduces the risk of overfitting. As a result, the model is encouraged to generalize better to unseen data, leading to improved performance during inference. More recent approaches have extended beyond dropout to implement more biologically-inspired pruning methods that permanently remove weights or connections based on importance criteria [9, 11], potentially offering advantages in both computational efficiency and performance.
In this work, we introduce a novel regularization method that more closely emulates biological synaptic pruning within deep neural networks, with a particular focus on enhancing time series forecasting performance. Unlike standard dropout, which randomly and temporarily deactivates neurons during training, or pruning techniques aimed solely at post-training model compression, our method applies an activity-dependent, gradual elimination of connections based on their contribution to model performance [12]. By combining temporal scheduling with context-sensitive pruning, our approach progressively weakens and permanently removes low-importance connections during training, enabling the network to adapt dynamically. By integrating directly into the training process, our method requires no additional pruning or fine-tuning stages and is especially effective in recurrent and temporal models commonly used for time series forecasting.
@@ 2 Related Works
In computational neuroscience and deep learning, several methods have been proposed to emulate biological synaptic pruning. LeCun et al. proposed Optimal Brain Damage [7], one of the earliest approaches to biologically inspired pruning in artificial neural networks. This method uses second-order derivatives of the loss function to estimate the saliency of each weight, pruning those with minimal impact on performance. However, this approach requires complete model training before pruning begins, making it a post-training optimization rather than an integrated learning process that could adapt to evolving temporal patterns in time series data.
@@
Han et al. [9] introduced magnitude-based pruning, where neural network weights with magnitudes below a certain threshold are eliminated, under the assumption that they contribute less to the final output. While effective in reducing network complexity, this approach primarily focuses on static weight values and does not consider the dynamic importance of individual nodes or features over time, which is particularly limiting for time series forecasting where temporal dependencies change throughout the learning process. This approach further requires three independent steps consisting of initial model training, pruning, and finally model fine-tuning, making it less practical for applications where continuous adaptation during training is valuable.
@@
Li et al. [13] proposed sensitivity-based pruning for Convolutional Networks, also known as contribution-aware pruning, which evaluates the increase in loss when individual neurons are ablated, thereby preserving units that significantly impact performance. This method aligns with the biological concept of activity-dependent synaptic retention, where synapses that contribute more to network function are retained, while those with minimal impact are pruned. However, this approach still operates through discrete evaluation phases rather than continuous monitoring, which may miss the evolving importance of features during model training and inference.
@@
Gradient-based pruning was proposed by Molchanov et al. [11], which analyzes the gradient flow through each neuron or connection and prunes those with consistently low gradient magnitudes, indicating minimal contribution to loss reduction. While this approach captures some dynamic behavior of the network, it often incurs high computational overhead due to the need to calculate gradients for every neuron or connection during training, and typically requires separate pruning phases distinct from regular training.
@@
Hassabis et al. [8] presented a broader framework linking neuroscience and artificial intelligence, emphasizing how mechanisms such as synaptic pruning and activity-dependent plasticity inform efficient computation in the brain. They argued that integrating biological principles into artificial systems, such as pruning underused or redundant connections, can lead to more robust, adaptable, and generalizable models. This perspective supports pruning not merely as a compression technique, but as a fundamental principle for learning and generalization, grounded in neuro-biological theory, which is particularly relevant for time series forecasting where adaptability to changing patterns is crucial.
@@
Frankle et al. [12] introduced the Lottery Ticket Hypothesis (LTH), which suggests that within large networks exist sparse sub-networks, referred to as ”winning tickets,” that can achieve comparable performance if identified early and pruned effectively. LTH finds inherently trainable sparse architectures that require retraining from specific initial conditions. This approach mimics developmental pruning by identifying and retaining strong sub-networks early in training, much like how the brain strengthens key neural connections during development while pruning others. However, the iterative process of training, pruning, and retraining may be less practical for applications where continuous adaptation during a single training cycle is preferred.
@@
Gradient-based Single-shot Pruning (SNIP) was introduced by Lee et al. [14], which estimates the importance of each connection at initialization by measuring the sensitivity of the loss to small perturbations in the weights. Connections with low gradient-based saliency scores are pruned before training begins, under the assumption that their contribution to learning will remain negligible. While computationally efficient, this approach makes one-time pruning decisions based on the network’s initial state, assuming that early indicators reliably predict long-term relevance, which may not hold for cases where feature importance can evolve as the model learns different temporal patterns.
@@
Wang et al. [15] proposed Gradient Signal Preservation (GraSP), aiming to maintain train-ability in pruned networks by preserving the gradient flow at initialization. GraSP evaluates the influence of each weight on the network’s Hessian-gradient product and prunes connections that minimally affect gradient propagation. This method improves over naive pruning by considering second-order interactions. However, like SNIP, it makes static pruning decisions at initialization rather than adapting to changing feature importance throughout training, which limits its effectiveness when temporal relationships may shift during the learning process.
@@ 3.1 Proposed Pruning Method
The method proposed in this study implements a global magnitude-based weight pruning strategy that progressively removes connections based on their importance to the network’s predictive performance. Our synaptic pruning approach differs fundamentally from standard dropout [10] and other regularization techniques in several key aspects. Specifically, our method:
@@
Our proposed method operates at the individual weight level using global magnitude comparison, ensuring that only the least important connections across the entire network are removed. This approach aligns well with biological synaptic pruning principles where weaker synaptic connections are eliminated to optimize neural efficiency [16, 17, 18].
@@ 3.4 Biological Inspiration
This approach mirrors biological synaptic pruning processes observed in neural development, where initially overproduced synaptic connections are selectively eliminated based on their functional importance [19, 16, 20]. The gradual nature of our cubic scheduling function reflects the progressive refinement seen in biological neural networks, where pruning occurs over extended periods rather than through abrupt elimination. Additionally, the global magnitude-based selection mechanism ensures that only the most functionally relevant connections are preserved, similar to how biological systems maintain synapses that contribute most effectively to information processing and learning.
@@ 3.5 Comparison with Alternative Approaches
Unlike alternative dropout approaches [21, 22], which introduce learnable parameters and stochastic behavior, our method provides deterministic, magnitude-based pruning that results in genuinely sparse networks. This approach differs from methods like SNIP [14] by operating during training rather than at initialization, allowing for dynamic adaptation to learned representations. The global comparison strategy ensures more balanced pruning across layers compared to layer-wise approaches, leading to more efficient overall network architectures.
@@ 3.7 Datasets
Two financial datasets were incorporated: a single stock ticker from the S&P 500 Stocks dataset [23] containing 1,259 records with 20 features, and the Bitcoin Daily historical dataset [24] with 365 records and 7 features. Both financial datasets are classified as low complexity, representing well-structured time series data with established patterns typical of financial markets.
The Household Electric Power Consumption dataset [25] provides the largest sample size with over 2 million records (2,075,259) and 9 features, offering extensive temporal data for energy usage analysis. Despite its substantial size, it maintains low complexity due to the straightforward nature of power consumption measurements.
Finally, the Air Quality dataset [26] presents a moderate-sized dataset with 9,446 records but the highest feature dimensionality at 313 features, resulting in high complexity. This dataset challenges our proposed methodology with high-dimensional environmental data encompassing multiple pollutants and atmospheric conditions. The diversity in dataset characteristics, ranging from 365 to over 2 million records, 7 to 313 features, and varying complexity levels, ensures robust evaluation of the proposed method across different data scenarios and computational challenges.
@@
S&P 500 Stocks [23] 1259 20 Low
Bitcoin Daily [24] 365 7 Low
Household Electric Power Consumption [25] 2075259 9 Low
Air Quality [26] 9446 313 High
@@ 3.8 Experimentation
A total of 96 experiments were performed using RNN, LSTM, and PatchTST architectures, measuring Mean Absolute Error (MAE) when using no regularization, standard dropout, Monte Carlo dropout [27] and our proposed pruning method, using input sequence lengths of 1, 3, 7, 14, 30 and 60. Each experiment was run for 10 trials using varying random seed initialization across a maximum of 20 epochs per trial, with the means of MAE and 95% Confidence Interval (CI) calculated across trials. Experimentation was performed using an Intel i9 20-core CPU and a single Nvidia RTX2080ti GPU. The Python programming language version 3.13.2 with PyTorch version 2.6 was used for all experimentation, with the full source code and utilized data publicly hosted at https://github.com/xalentis/SynapticPruning.
@@ 4.4 Computational Efficiency and Network Sparsification
The proposed synaptic pruning method demonstrates a positive trade-off between computational efficiency and performance across varying sequence lengths. The runtime performance comparison (Figure 6) reveals that while synaptic pruning initially exhibits higher computational overhead at short sequences, it achieves efficiency gains at longer sequences. At sequence length 7, synaptic pruning reaches its optimal performance, significantly outperforming both standard dropout and Monte Carlo dropout (which shows gradual degradation). However, the method exhibits a characteristic U-shaped performance curve, with runtime increasing substantially for very long sequences, though still maintaining competitive performance compared to the baseline without dropout.
The memory usage analysis (Figure 7) provides insights into the method’s scalability characteristics. All approaches demonstrate similar memory consumption patterns, with peak usage stabilizing after sequence length 3. This memory consistency across different sequence lengths suggests that the pruning mechanism effectively manages computational resources without introducing significant memory overhead.
Figure 6: Average pruning sparsity rates by model.
Figure 7: Average memory reduction rates by model.
@@ 4.5 Biological Plausibility and Neuro-scientific Validation
Our proposed method’s biological inspiration extends beyond superficial analogies to incorporate quantifiable aspects of synaptic pruning mechanisms, mirroring activity-dependent synaptic strengthening observed in long-term potentiation (LTP) [28] and long-term depression (LTD) [29]. The temporal dynamics of our pruning process align with developmental timescales observed in cortical maturation [30]. This biological grounding distinguishes our approach from purely engineering-based pruning methods and suggests potential for further bio-inspired enhancements.
Performance results reveal that synaptic pruning is particularly well-suited for moderate sequence lengths (3-14), where it achieves optimal efficiency. This behavior pattern aligns with the adaptive nature of biological synaptic pruning, where neural circuits optimize themselves for frequently encountered input patterns. The method’s ability to maintain stable memory usage while providing variable runtime performance could be particularly valuable for applications with diverse sequence length requirements [31, 32]. Our analysis shows that networks converge to optimal sparsity levels within the first 15 epochs, after which the pruning rate stabilizes. This early convergence behavior aligns with biological observations of critical periods of brain development [33, 34], suggesting that our method captures fundamental principles of neural circuit optimization during these developmental windows.

### v2

@@ Abstract
Biological synaptic pruning removes weak neural connections to improve efficiency, while standard dropout in artificial networks randomly deactivates neurons without considering connection importance. We propose a magnitude-based synaptic pruning method that better emulates biological processes by gradually removing connections according to their contribution to model performance. Integrated directly into the training loop as a dropout replacement, our method computes weight importance from absolute magnitudes across layers and applies a cubic schedule to progressively increase global sparsity. At regular intervals, pruning masks are updated by thresholding weights, permanently removing low-importance connections while preserving gradient flow for active ones. This continuous, data-driven pruning removes the need for separate pruning and fine-tuning phases. We evaluated the method across multiple time series forecasting architectures, including Recurrent Neural Networks, Long Short-Term Memory, and Patch Time Series Transformer models, using four datasets. Our synaptic pruning approach achieved the best overall performance ranking across all architectures, with statistically significant improvements confirmed by Friedman tests ($p<0.01$). In financial forecasting tasks, it reduced Mean Absolute Error by up to 20% compared to models using no dropout or standard dropout, with reductions reaching 52% in select transformer models. The proposed mechanism advances regularization by coupling dynamic weight elimination with progressive sparsification during training. Its ease of integration and consistent accuracy gains make it a practical alternative to dropout, particularly for high-precision domains such as financial time series forecasting. Source code is available at https://github.com/xalentis/SynapticPruning
@@
journal: Applied Soft Computing
@@ 1 Introduction
Synaptic pruning is a fundamental neuro-developmental process whereby weak or redundant synaptic connections are selectively eliminated to enhance the efficiency of neuronal networks rakic1986 . This process is highly activity-dependent, meaning that synapses with lower activity or utility are more likely to be pruned, resulting in improved energy efficiency and functional specialization Huttenlocher1997 . During critical periods of development, the human brain initially overproduces synapses, creating a dense network of connections that undergoes substantial refinement through adolescence and early adulthood Chechik1998 , Selemon1999 . This refinement is essential for optimal cognitive function and normal brain development, with disruptions in pruning mechanisms associated with various neuro-developmental disorders Zhan2014 , Tang2014 .
In the field of artificial neural networks, researchers have drawn inspiration from biological synaptic pruning to develop methods for improving network efficiency and generalization capacity LeCun1990 , Hassabis2017 , Han2015 . Dropout regularization Srivastava2014 functions by randomly deactivating, or ”dropping out,” a subset of neurons during each training iteration. This process forces the network to avoid over-reliance on any particular neuron, promoting redundancy and robustness in the learned representations. By preventing neurons from co-adapting too strongly to specific patterns in the training data, dropout reduces the risk of overfitting. As a result, the model is encouraged to generalize better to unseen data, leading to improved performance during inference. More recent approaches have extended beyond dropout to implement more biologically-inspired pruning methods that permanently remove weights or connections based on importance criteria Han2015 , Molchanov2017 , potentially offering advantages in both computational efficiency and performance.
In this work, we introduce a novel regularization method that more closely emulates biological synaptic pruning within deep neural networks, with a particular focus on enhancing time series forecasting performance. Unlike standard dropout, which randomly and temporarily deactivates neurons during training, or pruning techniques aimed solely at post-training model compression, our method applies an activity-dependent, gradual elimination of connections based on their contribution to model performance Frankle2019 . By combining temporal scheduling with context-sensitive pruning, our approach progressively weakens and permanently removes low-importance connections during training, enabling the network to adapt dynamically. By integrating directly into the training process, our method requires no additional pruning or fine-tuning stages and is especially effective in recurrent and temporal models commonly used for time series forecasting.
@@ 2 Related Works
In computational neuroscience and deep learning, several methods have been proposed to emulate biological synaptic pruning. LeCun et al. proposed Optimal Brain Damage LeCun1990 , one of the earliest approaches to biologically inspired pruning in artificial neural networks. This method uses second-order derivatives of the loss function to estimate the saliency of each weight, pruning those with minimal impact on performance. However, this approach requires complete model training before pruning begins, making it a post-training optimization rather than an integrated learning process that could adapt to evolving temporal patterns in time series data.
@@
Han et al. Han2015 introduced magnitude-based pruning, where neural network weights with magnitudes below a certain threshold are eliminated, under the assumption that they contribute less to the final output. While effective in reducing network complexity, this approach primarily focuses on static weight values and does not consider the dynamic importance of individual nodes or features over time, which is particularly limiting for time series forecasting where temporal dependencies change throughout the learning process. This approach further requires three independent steps consisting of initial model training, pruning, and finally model fine-tuning, making it less practical for applications where continuous adaptation during training is valuable.
@@
Li et al. Li2016 proposed sensitivity-based pruning for Convolutional Networks, also known as contribution-aware pruning, which evaluates the increase in loss when individual neurons are ablated, thereby preserving units that significantly impact performance. This method aligns with the biological concept of activity-dependent synaptic retention, where synapses that contribute more to network function are retained, while those with minimal impact are pruned. However, this approach still operates through discrete evaluation phases rather than continuous monitoring, which may miss the evolving importance of features during model training and inference.
@@
Gradient-based pruning was proposed by Molchanov et al. Molchanov2017 , which analyzes the gradient flow through each neuron or connection and prunes those with consistently low gradient magnitudes, indicating minimal contribution to loss reduction. While this approach captures some dynamic behavior of the network, it often incurs high computational overhead due to the need to calculate gradients for every neuron or connection during training, and typically requires separate pruning phases distinct from regular training.
@@
Hassabis et al. Hassabis2017 presented a broader framework linking neuroscience and artificial intelligence, emphasizing how mechanisms such as synaptic pruning and activity-dependent plasticity inform efficient computation in the brain. They argued that integrating biological principles into artificial systems, such as pruning underused or redundant connections, can lead to more robust, adaptable, and generalizable models. This perspective supports pruning not merely as a compression technique, but as a fundamental principle for learning and generalization, grounded in neuro-biological theory, which is particularly relevant for time series forecasting where adaptability to changing patterns is crucial.
@@
Frankle et al. Frankle2019 introduced the Lottery Ticket Hypothesis (LTH), which suggests that within large networks exist sparse sub-networks, referred to as ”winning tickets,” that can achieve comparable performance if identified early and pruned effectively. LTH finds inherently trainable sparse architectures that require retraining from specific initial conditions. This approach mimics developmental pruning by identifying and retaining strong sub-networks early in training, much like how the brain strengthens key neural connections during development while pruning others. However, the iterative process of training, pruning, and retraining may be less practical for applications where continuous adaptation during a single training cycle is preferred.
@@
Gradient-based Single-shot Pruning (SNIP) was introduced by Lee et al. lee2019 , which estimates the importance of each connection at initialization by measuring the sensitivity of the loss to small perturbations in the weights. Connections with low gradient-based saliency scores are pruned before training begins, under the assumption that their contribution to learning will remain negligible. While computationally efficient, this approach makes one-time pruning decisions based on the network’s initial state, assuming that early indicators reliably predict long-term relevance, which may not hold for cases where feature importance can evolve as the model learns different temporal patterns.
@@
Wang et al. wang2020 proposed Gradient Signal Preservation (GraSP), aiming to maintain train-ability in pruned networks by preserving the gradient flow at initialization. GraSP evaluates the influence of each weight on the network’s Hessian-gradient product and prunes connections that minimally affect gradient propagation. This method improves over naive pruning by considering second-order interactions. However, like SNIP, it makes static pruning decisions at initialization rather than adapting to changing feature importance throughout training, which limits its effectiveness when temporal relationships may shift during the learning process.
@@ 3.1 Proposed Pruning Method
The method proposed in this study implements a global magnitude-based weight pruning strategy that progressively removes connections based on their importance to the network’s predictive performance. Our synaptic pruning approach differs fundamentally from standard dropout Srivastava2014 and other regularization techniques in several key aspects. Specifically, our method:
@@
Our proposed method operates at the individual weight level using global magnitude comparison, ensuring that only the least important connections across the entire network are removed. This approach aligns well with biological synaptic pruning principles where weaker synaptic connections are eliminated to optimize neural efficiency chechik1999 , lecun2021 , zhu2017 .
@@ 3.4 Biological Inspiration
This approach mirrors biological synaptic pruning processes observed in neural development, where initially overproduced synaptic connections are selectively eliminated based on their functional importance zhu2017b , chechik1999 , chklovskii2004 . The gradual nature of our cubic scheduling function reflects the progressive refinement seen in biological neural networks, where pruning occurs over extended periods rather than through abrupt elimination. Additionally, the global magnitude-based selection mechanism ensures that only the most functionally relevant connections are preserved, similar to how biological systems maintain synapses that contribute most effectively to information processing and learning.
@@ 3.5 Comparison with Alternative Approaches
Unlike alternative dropout approaches kingma2015 , gal2017 , which introduce learnable parameters and stochastic behavior, our method provides deterministic, magnitude-based pruning that results in genuinely sparse networks. This approach differs from methods like SNIP lee2019 by operating during training rather than at initialization, allowing for dynamic adaptation to learned representations. The global comparison strategy ensures more balanced pruning across layers compared to layer-wise approaches, leading to more efficient overall network architectures.
@@ 3.7 Datasets
Two financial datasets were incorporated: a single stock ticker from the S&P 500 Stocks dataset dataset_sp500 containing 1,259 records with 20 features, and the Bitcoin Daily historical dataset dataset_bitcoin with 365 records and 7 features. Both financial datasets are classified as low complexity, representing well-structured time series data with established patterns typical of financial markets.
The Household Electric Power Consumption dataset dataset_power provides the largest sample size with over 2 million records (2,075,259) and 9 features, offering extensive temporal data for energy usage analysis. Despite its substantial size, it maintains low complexity due to the straightforward nature of power consumption measurements.
Finally, the Air Quality dataset dataset_air presents a moderate-sized dataset with 9,446 records but the highest feature dimensionality at 313 features, resulting in high complexity. This dataset challenges our proposed methodology with high-dimensional environmental data encompassing multiple pollutants and atmospheric conditions. The diversity in dataset characteristics, ranging from 365 to over 2 million records, 7 to 313 features, and varying complexity levels, ensures robust evaluation of the proposed method across different data scenarios and computational challenges.
@@
S&P 500 Stocks dataset_sp500 1259 20 Low
Bitcoin Daily dataset_bitcoin 365 7 Low
Household Electric Power Consumption dataset_power 2075259 9 Low
Air Quality dataset_air 9446 313 High
@@ 3.8 Experimentation
A total of 96 experiments were performed using RNN, LSTM, and PatchTST architectures, measuring Mean Absolute Error (MAE) when using no regularization, standard dropout, Monte Carlo dropout gal2016 and our proposed pruning method, using input sequence lengths of 1, 3, 7, 14, 30 and 60. Each experiment was run for 10 trials using varying random seed initialization across a maximum of 20 epochs per trial, with the means of MAE and 95% Confidence Interval (CI) calculated across trials. Experimentation was performed using an Intel i9 20-core CPU and a single Nvidia RTX2080ti GPU. The Python programming language version 3.13.2 with PyTorch version 2.6 was used for all experimentation, with the full source code and utilized data publicly hosted at https://github.com/xalentis/SynapticPruning.
@@ 4.4 Computational Efficiency and Network Sparsification
The proposed synaptic pruning method demonstrates a positive trade-off between computational efficiency and performance across varying sequence lengths. The runtime performance comparison (Figure 6) reveals that while synaptic pruning initially exhibits higher computational overhead at short sequences, it achieves efficiency gains at longer sequences. At sequence lengths 1 and 60, synaptic pruning reaches its optimal performance (Figure 5), significantly outperforming both standard dropout and Monte Carlo dropout (which shows gradual degradation). However, the method exhibits an inverted U-shaped performance curve, though still maintaining competitive performance compared to the baseline without dropout. Where computational costs do increase, these are justified by substantial improvements in predictive accuracy, with the method consistently achieving lower error rates than baseline approaches, demonstrating that the additional computational investment delivers meaningful performance returns.
Figure 6: Average peak memory usage vs sequence length by model”.
@@ 4.5 Biological Plausibility and Neuro-scientific Validation
Our proposed method’s biological inspiration extends beyond superficial analogies to incorporate quantifiable aspects of synaptic pruning mechanisms, mirroring activity-dependent synaptic strengthening observed in long-term potentiation (LTP) bliss1993 and long-term depression (LTD) bear1996 . The temporal dynamics of our pruning process align with developmental timescales observed in cortical maturation bourgeois1997 . This biological grounding distinguishes our approach from purely engineering-based pruning methods and suggests potential for further bio-inspired enhancements.
Performance results reveal that synaptic pruning is particularly well-suited for moderate sequence lengths (3-14), where it achieves optimal efficiency. This behavior pattern aligns with the adaptive nature of biological synaptic pruning, where neural circuits optimize themselves for frequently encountered input patterns. The method’s ability to maintain stable memory usage while providing variable runtime performance could be particularly valuable for applications with diverse sequence length requirements howard2017 , iandola2016 . Our analysis shows that networks converge to optimal sparsity levels within the first 15 epochs, after which the pruning rate stabilizes. This early convergence behavior aligns with biological observations of critical periods of brain development hensch2004 , takesian2013 , suggesting that our method captures fundamental principles of neural circuit optimization during these developmental windows.